A solution of benzene in methanol has a transmittance of 83% in a 1.00 cm cell at a wavelength of 254 nm. Only the benzene absorbs light at this wavelength, not the methanol. What will the solution’s transmittance be if it is placed in a 10.00 cm long pathlength cell?

  • A. 2.1%
  • B. 27.8%
  • C. 15.5%
  • D. 96.2%
  • E. 98.2%
  • F. 8.80%
  • G. 8.30%
Answer #1

Answer is (c) 15.5%

As we know through Beer-Lambert's Law :

log(I0 / I)=constant*(length of solution the light passes through)*(concentration of solution) [keeping in mind the correct units though not required here)

we know through given conditions :

log(0.83)=constant*l*c

Therefore, constant*l*c=0.08

Now change is done only in concentration which is 10 times of original so,

Now , log(I0 / I)=10*constant*l*c

from here you will get the ratio and on multiplying with 100 we will get percentage.